# Language Models for Controllable DNA Sequence Design

### v1

@@ 1 Introduction
DNA sequence design is a transformative scientific endeavor that revolutionizes our understanding of biology and catalyzes major advances across healthcare, agriculture, and environmental conservation, etc (Zrimec et al., 2022; Killoran et al., 2017). In these tasks, it is commonly expected that the generated DNA sequence achieves some biological outcomes (Uehara et al., 2025; Li et al., 2025). Thus, controllable generation is of practical significance (Nguyen et al., 2024a) in which the sequence generation is guided by various properties, e.g., binding to specific proteins or exhibiting particular transcription activation likelihoods.
@@
Recent advances in generative models have triggered widespread interests in DNA sequence design using diffusion (Avdeyev et al., 2023; Li et al., 2024; Sarkar et al., 2024) and flow matching (Stark et al., 2024) generative methods. These methods have shown promise, particularly in modeling global structure and optimizing over continuous latent spaces. However, they are not inherently tailored to discrete, symbolic sequence generation—properties that are central to DNA sequences. While those methods are in general effective in generation task, they are not naturally designed for generating discrete and variable-length sequences. In contrast, Language models (LMs) offer an alternative perspective: they are naturally suited for discrete, variable-length generation and have achieved remarkable success in analogous domains like natural language. In this work, we explore the potential of LMs for controllable DNA sequence design, proposing them as a complementary approach to existing generative methods.
@@
In this paper, we aim to explore the use of transformer-based language models for DNA sequence design conditioned on specific biological properties (Vaswani et al., 2017; Stark et al., 2024; Sarkar et al., 2024; Su et al., 2025). To this end, we develop ATGC-Gen, an Automated Transformer Generator for Controllable Generation. We instantiate ATGC-Gen with both decoder-only (e.g., GPT) and encoder-only (e.g., BERT) transformer architectures to examine their respective capabilities for controllable generation. It is designed to encode and integrate heterogeneous biological information, such as cell types, protein sequences, and transcription activation signals, into the sequence generation process. This unified framework enables ATGC-Gen to capture complex relationships between DNA sequences and their biological contexts, facilitating the generation of sequences with specific desirable properties.
@@
To evaluate the effectiveness of our proposed approach, we apply ATGC-Gen, a transformer-based language model that incorporates cross-modal biological information for controllable DNA sequence generation. By conditioning on diverse properties such as cell types, transcription factors, and regulatory signals, ATGC-Gen enables flexible and biologically grounded sequence design. We first assess its performance on established promoter (Avdeyev et al., 2023; Stark et al., 2024; Sarkar et al., 2024) and enhancer (Taskiran et al., 2024) generation tasks. To further explore its capability in handling complex biological contexts, we introduce a new dataset based on ChIP-Seq experiments, which involves generating sequences that bind to specific proteins in specific cell types. Experimental results across different tasks, evaluated in terms of functionality, fluency, and diversity, demonstrate that ATGC-Gen achieves strong and consistent performance, highlighting the promise of language models for advancing controllable and property-aware DNA sequence design.
@@ 2 Background and Related Work
DNA Generation. We study the problem of DNA sequence generation. In the field of synthetic biology and genetic engineering, unconditional or random DNA generation has limited applications, since DNA sequences are designed or edited to achieve specific outcomes, e.g., producing a protein or altering a metabolic pathway. An unconditionally generated DNA sequence lacks the necessary context to be meaningful or useful in biological systems. Hence, in this work, we focus on studying the task of DNA generation given specific properties, such as a particular organism, cell type, or functional requirement. Formally, we define a DNA dataset as $\mathcal{U}=\{(U_{1},C_{1}),(U_{2},C_{2}),\cdots,(U_{N},C_{N})\}$, where $U_{i}$ is a DNA sequence and $C_{i}$ is its corresponding biological property. We aim to learn a conditional generative model $p_{\theta}(\cdot|C_{i})$ on $\mathcal{U}$ so that the model $G$, parameterized with $\theta$, can generate DNA sequences fulfilling $C_{i}$. Previous studies have made several initial attempts to apply deep learning methods for controllable DNA generation. The Dirichlet Diffusion Score Model (DDSM) (Avdeyev et al., 2023) uses Dirichlet distribution to discretize the diffusion process, effectively generating DNA sequences. In Stark et al. (2024) the Dirichlet distribution is used in the flow matching process, enhancing the quality of generated DNA sequences. DiscDiff (Li et al., 2024) employs Latent Discrete Diffusion model on the DNA generation task and proposes a post-training refinement algorithm to improve the generation quality. D3 (Sarkar et al., 2024) uses score entropy for discrete diffusion on the conditional DNA sequence generation.
DNA Language Model. Recent research has demonstrated the power of language models for learning from discrete sequences, which is ideally suited for modeling DNA sequences. A few recent studies have employed language models for encoding DNA sequences in prediction tasks. Early works adopt Transformers to encode DNA sequence data, and notable examples include DNABERT, DNABERT-2 and Nucleotide Transformer (Ji et al., 2021; Zhou et al., 2023b; Dalla-Torre et al., 2023). Recent works focus more on using state-space models, producing models like HyenaDNA (Nguyen et al., 2024b) and Mamba (Gu & Dao, 2023). HyenaDNA employs the Hyena operator (Poli et al., 2023) and uses implicit convolution to handle long DNA sequences up to one million bases effectively. Mamba captures long-range interactions in a simplified manner, demonstrating significant effectiveness in long sequence modeling (Gu & Dao, 2023; Schiff et al., 2024). Caduceus (Schiff et al., 2024) improves Mamba based on bi-directional encoding and reverse complement nature of DNA sequence. Both HyenaDNA and Mamba have been pretrained using the next token prediction paradigm, making them particularly well-suited for DNA generation tasks. Based on the state space model, Evo (Nguyen et al., 2024a) publishes a pretrained model encoding DNA, RNA, protein modalities, demonstrating impressive performance across a variety of tasks in these modalities. However, current studies invariably focus on DNA prediction tasks and do not involve generation. On the other hand, the true power of language models lies in their generation capabilities. Hence, our primary aim in this work is to apply language models for controllable DNA generation, generating sequences to satisfy predefined biological properties across various modalities.
@@ 3.1 Representation Encoding
Let the DNA embedding sequence be $\mathbf{H}_{e}=\{\mathbf{h}_{e,1},\cdots,\mathbf{h}_{e,n}\}\in\mathbb{R}^{n\times d_{h}}$. We prepend the property embedding to the DNA sequence to form the final transformer input:
@@
$\tilde{\mathbf{x}}_{i}=\text{concat}(\mathbf{x}_{e,i},\mathbf{s}_{i})\in\mathbb{R}^{4+d_{s}}.$
@@
$\tilde{\mathbf{h}}_{e,i}=W_{f}\tilde{\mathbf{x}}_{i}+\mathbf{b}_{f},$
@@ 3.2 Training Objectives
In this setting, we train a decoder-only language model $F_{\theta}$ to generate DNA sequences in an autoregressive way, conditioned on the property representations $\mathbf{H}_{{p}}$. The input sequence consists of the property tokens followed by the DNA tokens, and the model is trained to predict the next nucleotide at each position. This training paradigm naturally complements sequence-level integration, where the prepended property tokens act as a global prefix that controls all subsequent predictions.
Following standard practice (Zhou et al., 2023b; Schiff et al., 2024; Radford et al., 2018), we use a single-nucleotide tokenizer, where each base (A, C, G, T) is treated as a discrete token. The model is trained using the next-token prediction objective with cross-entropy loss:
$\min_{\theta}\,\,\mathbb{E}_{U\sim\mathcal{U}}\left[\sum_{i=1}^{|U|-1}\ell\left(F_{\theta}\left(\mathbf{H}_{p},u_{1},\cdots,u_{i}\right),u_{i+1}\right)\right],$ (3)
where $U=\{u_{1},\cdots,u_{|U|}\}$ is a DNA sequence from the dataset $\mathcal{U}$, and $\ell$ is the cross-entropy loss over the predicted nucleotide distribution.
@@
In parallel, we consider an encoder-based training approach using masked language modeling (MLM), similar to BERT (Devlin et al., 2019). This formulation requires sequences of fixed length and is trained to recover randomly masked tokens in the input. Specifically, given an input sequence of DNA tokens $U=\{u_{1},\cdots,u_{n}\}$ and property representation $\mathbf{H}_{p}$, we randomly select a subset of positions $\mathcal{M}\subset\{1,\cdots,n\}$ to mask, and train the model $F_{\theta}$ to reconstruct the original nucleotides at those positions:
@@ 3.3 Controllable Generation
To promote diversity and controllability, we sample tokens using temperature sampling (Ackley et al., 1985; Ficler & Goldberg, 2017).
@@
Remark. This formulation shares structural similarities with recent masked denoising approaches, such as masked discrete diffusion (Sahoo et al., 2024; Nie et al., 2025), where a corrupted input is progressively reconstructed through iterative masked token prediction.
@@ 4 Controllable DNA Generation with ChIP-Seq
We develop a dataset derived from ChIP-Seq experiments for evaluating controllable DNA generation problems. ChIP-Seq experiments aim to identify the binding sites of DNA-associated proteins on the genome, specifically the DNA sequences where proteins physically attach or bind. In our setting, we focus on predicting potential binding DNA sequences of variable lengths based on given protein sequences across different cell types. This prediction can advocate our understanding of the complex networks of gene regulation and aid in designing DNA sequences with specific functionalities for genetic engineering.
Our proposed dataset is obtained from a comprehensive collection of ChIP-Seq experiments generated by the ENCODE project (Consortium et al., 2012), which identifies specific DNA sequences to which proteins bind. We filter the raw data by considering the highest binding scores, length properties and other criteria, and detailed steps are provided in the Appendix A.
@@
The DNA sequences are from GRCh38 (HG38) reference human genome assembly, used for mapping and aligning genetic data. Each base pair in the DNA sequence is one of the four nucleotides: adenine (A), cytosine (C), guanine (G), thymine (T). Given the protein name, we can query protein databases (Berman et al., 2000) to find the corresponding protein sequences and structures.
@@
  • Functionality: This metric measures how well the generated DNA sequence binds with the given protein under a specific cell type. We use the SEI framework (Chen et al., 2022), which employs a deep learning model to predict 21,907 chromatin profiles based on DNA sequences. As shown in Figure 2, each chromatin profile is defined by cell type, protein name and experiment. The prediction value indicates the probability of binding between the DNA and the protein. We focus exclusively on chromatin profiles derived from ENCODE experiments, as our ChIP-Seq dataset is sourced from the ENCODE project.
  • Fluency: Similar to the fluency used in the text generation task (Li et al., 2022; Zhou et al., 2023a), we adopt it to assesses how smooth and natural the generated DNA sequence is. Fluency is defined as the perplexity of the generated DNA in a pretrained language model, with lower value indicating better fluency. We use the HyenaDNA pretrained model as the base model $F_{b}$ (Nguyen et al., 2024b). Given the cross-entropy loss $\ell$, the fluency of a single DNA sequence $U=\{u_{1},\cdots,u_{n}\}$ is the exponential of the loss
@@
  • Diversity: In a specific cell type, a given protein can bind to many different DNA sequences. This metric measures the diversity of the generated sequences. Following Li et al. (2024), the diversity score is defined as
@@ 5.1 Promoter Generation
Dataset and task descriptions. This dataset contains 100,000 promoter sequences from GRCh38 (HG38) human reference genome. Each sequence is annotated with the transcription initiation profile, indicating the likelihood of transcription initiation activity at each base pair (fan, 2014). We use the same DNA sequences as in previous works (Avdeyev et al., 2023; Stark et al., 2024), specifically splitting 1,024 base pair long DNA sequences around each annotated transcription start site position (Hon et al., 2017). The task is to interpret the transcription initiation profile into the corresponding DNA sequences of the same length.
Metrics. Following prior works (Avdeyev et al., 2023; Stark et al., 2024), we evaluate the generation performance using the mean squared error (MSE) between the predicted regulatory activity of the generated and original DNA sequences. The regulatory activity is computed using the SEI framework (Chen et al., 2022), a published deep learning model that predicts 21,907 regulatory features for a given DNA sequence. For our primary evaluation, we extract promoter-related activity by focusing on the H3K4me3 chromatin mark. In addition, we report the Kolmogorov–Smirnov (KS) statistic between the distributions of regulatory activity values from the generated and original sequences, as proposed in (Sarkar et al., 2024), to assess distributional alignment. We further evaluate the fluency of the generated sequences by HyenaDNA (Nguyen et al., 2023), as described in Section 4, to ensure that generated DNA is biologically plausible and structurally coherent.

### v2

@@ 1 Introduction
DNA sequence design is a transformative scientific endeavor that revolutionizes our understanding of biology and catalyzes major advances across healthcare, agriculture, and environmental conservation, etc (zrimec2022controlling; killoran2017generating). In these tasks, it is commonly expected that the generated DNA sequence achieves some biological outcomes (uehara2025reward; li2025dynamic). Thus, controllable generation is of practical significance (evo) in which the sequence generation is guided by various properties, e.g., binding to specific proteins or exhibiting particular transcription activation likelihoods.
@@
Recent advances in generative models have triggered widespread interests in DNA sequence design using diffusion (avdeyev2023dirichlet; discdiff; sarkar2024designing) and flow matching (stark2024dirichlet) generative methods. These methods have shown promise, particularly in modeling global structure and optimizing over continuous latent spaces. However, they are not inherently tailored to discrete, symbolic sequence generation—properties that are central to DNA sequences. While those methods are in general effective in generation task, they are not naturally designed for generating discrete and variable-length sequences. In contrast, Language models (LMs) offer an alternative perspective: they are naturally suited for discrete, variable-length generation and have achieved remarkable success in analogous domains like natural language. In this work, we explore the potential of LMs for controllable DNA sequence design, proposing them as a complementary approach to existing generative methods.
@@
In this paper, we aim to explore the use of transformer-based language models for DNA sequence design conditioned on specific biological properties (vaswani2017attention; stark2024dirichlet; sarkar2024designing; su2025learning). To this end, we develop ATGC-Gen, an Automated Transformer Generator for Controllable Generation. We instantiate ATGC-Gen with both decoder-only (e.g., GPT) and encoder-only (e.g., BERT) transformer architectures to examine their respective capabilities for controllable generation. It is designed to encode and integrate heterogeneous biological information, such as cell types, protein sequences, and transcription activation signals, into the sequence generation process. This unified framework enables ATGC-Gen to capture complex relationships between DNA sequences and their biological contexts, facilitating the generation of sequences with specific desirable properties.
@@
To evaluate the effectiveness of our proposed approach, we apply ATGC-Gen, a transformer-based language model that incorporates cross-modal biological information for controllable DNA sequence generation. By conditioning on diverse properties such as cell types, transcription factors, and regulatory signals, ATGC-Gen enables flexible and biologically grounded sequence design. We first assess its performance on established promoter (avdeyev2023dirichlet; stark2024dirichlet; sarkar2024designing) and enhancer (taskiran2024cell) generation tasks. To further explore its capability in handling complex biological contexts, we introduce a new dataset based on ChIP-Seq experiments, which involves generating sequences that bind to specific proteins in specific cell types. Experimental results across different tasks, evaluated in terms of functionality, fluency, and diversity, demonstrate that ATGC-Gen achieves strong and consistent performance, highlighting the promise of language models for advancing controllable and property-aware DNA sequence design.
@@ 2 Background and Related Work
DNA Generation. We study the problem of DNA sequence generation. In the field of synthetic biology and genetic engineering, unconditional or random DNA generation has limited applications, since DNA sequences are designed or edited to achieve specific outcomes, e.g., producing a protein or altering a metabolic pathway. An unconditionally generated DNA sequence lacks the necessary context to be meaningful or useful in biological systems. Hence, in this work, we focus on studying the task of DNA generation given specific properties, such as a particular organism, cell type, or functional requirement. Formally, we define a DNA dataset as $\mathcal{U}=\{(U_{1},S_{1}),(U_{2},S_{2}),\cdots,(U_{N},S_{N})\}$, where $U_{i}$ is a DNA sequence and $S_{i}$ is its corresponding biological property. We aim to learn a conditional generative model $p_{\theta}(\cdot|S_{i})$ on $\mathcal{U}$ so that the model $F_{\theta}$, parameterized with $\theta$, can generate DNA sequences fulfilling $S_{i}$. Previous studies have made several initial attempts to apply deep learning methods for controllable DNA generation. The Dirichlet Diffusion Score Model (DDSM) (avdeyev2023dirichlet) uses Dirichlet distribution to discretize the diffusion process, effectively generating DNA sequences. In stark2024dirichlet the Dirichlet distribution is used in the flow matching process, enhancing the quality of generated DNA sequences. DiscDiff (discdiff) employs Latent Discrete Diffusion model on the DNA generation task and proposes a post-training refinement algorithm to improve the generation quality. D3 (sarkar2024designing) uses score entropy for discrete diffusion on the conditional DNA sequence generation.
DNA Language Model. Recent research has demonstrated the power of language models for modeling biomolecular discrete sequences (wang2024lc; sgarbossa2025protmamba; xu2024smiles), which is ideally suited for modeling DNA sequences. A few recent studies have employed language models for encoding DNA sequences in prediction or representation learning tasks. Early works adopt Transformers to encode DNA sequence data, and notable examples include DNABERT, DNABERT-2 and Nucleotide Transformer (dnabert; dnabert2; NT). More recently, state space models such as HyenaDNA (hyenaDNA) and Mamba (gu2023mamba) have been applied to long DNA sequences, with Caduceus (schiff2024caduceus) further improving Mamba with bi-directional encoding and reverse complement awareness. Evo (evo) extends this line of work by pretraining on DNA, RNA, and protein modalities. In parallel, Bio-xLSTM (schmidinger2024bio) explores xLSTM architectures for biological sequences, which, similar to state space models, achieve a linear runtime dependency on sequence length and provide efficient sequence modeling. These approaches primarily propose efficient modeling architectures for biological sequences and have achieved strong results in prediction benchmarks. While powerful for encoding and prediction, these works do not directly address explicit controllable sequence design. In practice, although uni-modal models can achieve limited controllability by conditioning on sequences with desirable properties, generating condition-specific DNA sequences typically requires re-training or fine-tuning a separate model for each condition. By contrast, our framework integrates heterogeneous biological properties as conditioning inputs, providing a unified and scalable way to support diverse design criteria.
Cross-Modality Representation Learning in Biology. Many works combine heterogeneous modalities in latent spaces in biomolecular modeling. For example, some works (schimunek2023context; fifty2023context; zhou2025multimodal) fuse molecular structure with other modality information, achieving strong performance on molecular property prediction benchmarks. Other efforts integrate different modalities of protein information to improve representation learning (gligorijevic2021structure; brandes2022proteinbert; hayes2025simulating). However, these approaches focus primarily on prediction tasks rather than sequence generation, and the modalities considered are typically molecular or protein structure and activity labels. In contrast, our work extends cross-modal conditioning to the setting of controllable DNA sequence generation, where diverse biological attributes are incorporated as conditioning signals to guide the design of functional DNA sequences.
@@ 3.1 Representation Encoding
Let the DNA embedding sequence be $\mathbf{H}_{e}=\{\mathbf{h}_{e,1},\cdots,\mathbf{h}_{e,n}\}\in\mathbb{R}^{n\times d_{h}}$, where each $\mathbf{h}_{e,i}$ is obtained by mapping the $i$-th DNA token through a learnable embedding layer that projects discrete nucleotides into a $d_{h}$-dimensional continuous space. We prepend the property embedding to the DNA sequence to form the final transformer input:
@@
$\tilde{\mathbf{x}}_{i}=\text{concat}(\mathbf{x}_{e,i},\mathbf{s}_{i})\in\mathbb{R}^{(4+d_{s})\times 1}.$
@@
$\tilde{\mathbf{h}}_{e,i}=W_{f}\tilde{\mathbf{x}}_{i}+\mathbf{b}_{f},$ (2)
@@ 3.2 Training Objectives
In this setting, we train a decoder-only language model $F_{\theta}$ to generate DNA sequences autoregressively. To unify notation, let $X_{i}$ denote the model input at step $i$: (i) under sequence-level integration, $X_{i}=\{\mathbf{H}_{p};u_{1},\ldots,u_{i}\}$, where the prepended property tokens act as a global prefix; (ii) under feature-level integration, $X_{i}=\{\tilde{\mathbf{h}}_{e,1},\ldots,\tilde{\mathbf{h}}_{e,i}\}$, where $\tilde{\mathbf{h}}_{e,j}$ is defined in equation 2.
Following standard practice (dnabert2; schiff2024caduceus; radford2018improving), we use a single-nucleotide tokenizer, where each base (A, C, G, T) is treated as a discrete token. The model is trained using the next-token prediction objective with cross-entropy loss
$\min_{\theta}\,\,\mathbb{E}_{U\sim\mathcal{U}}\left[\sum_{i=1}^{|U|-1}\ell\big(F_{\theta}(X_{i}),\,u_{i+1}\big)\right],$ (4)
where $U=\{u_{1},\cdots,u_{|U|}\}$ is a DNA sequence from the dataset $\mathcal{U}$ and $\ell$ is the cross-entropy loss over the predicted nucleotide distribution.
@@
In parallel, we consider an encoder-based training approach using masked language modeling (MLM), similar to BERT (devlin2019bert). This formulation requires sequences of fixed length and is trained to recover randomly masked tokens in the input. Specifically, given an input embedding sequence $\mathbf{H}_{\text{input}}$ (constructed from both property and DNA tokens as described in Section 3.1), we randomly select a subset of positions $\mathcal{M}\subset\{1,\cdots,n\}$ to mask, and train the model $F_{\theta}$ to reconstruct the original nucleotides at those positions:
@@ 3.3 Controllable Generation
To promote diversity and controllability, we sample tokens using temperature sampling (ackley1985learning; ficler2017controlling).
@@
Remark. This formulation shares structural similarities with recent masked denoising approaches, such as masked discrete diffusion (sahoo2024simple; nie2025large), where a corrupted input is progressively reconstructed through iterative masked token prediction.
@@ 4 Controllable DNA Generation with ChIP-Seq
While existing DNA datasets are sufficient to benchmark the DNA generation tasks, they do not capture the variable-length nature of many real biological sequences. We therefore curate a variable-length dataset to highlight the adaptability of our framework to more realistic and flexible generation scenarios. Therefore, we develop a dataset derived from ChIP-Seq experiments for evaluating controllable DNA generation problems. ChIP-Seq experiments aim to identify the binding sites of DNA-associated proteins on the genome, specifically the DNA sequences where proteins physically attach or bind. In our setting, we focus on predicting potential binding DNA sequences of variable lengths based on given protein sequences across different cell types. This prediction can advocate our understanding of the complex networks of gene regulation and aid in designing DNA sequences with specific functionalities for genetic engineering.
Our proposed dataset is obtained from a comprehensive collection of ChIP-Seq experiments generated by the ENCODE project (encode2012integrated), which identifies specific DNA sequences to which proteins bind. We filter the raw data by considering the highest binding scores, length properties and other criteria, and detailed steps are provided in the Appendix A.
@@
The DNA sequences are from GRCh38 (HG38) reference human genome assembly, used for mapping and aligning genetic data. Each base pair in the DNA sequence is one of the four nucleotides: adenine (A), cytosine (C), guanine (G), thymine (T). Given the protein name, we can query protein databases (PDB) to find the corresponding protein sequences and structures.
@@
  • Functionality: This metric measures how well the generated DNA sequence binds with the given protein under a specific cell type. We use the SEI framework (SEI), which employs a deep learning model to predict 21,907 chromatin profiles based on DNA sequences. As shown in Figure 2, each chromatin profile is defined by cell type, protein name and experiment. The prediction value indicates the probability of binding between the DNA and the protein. We focus exclusively on chromatin profiles derived from ENCODE experiments, as our ChIP-Seq dataset is sourced from the ENCODE project.
  • Fluency: Similar to the fluency used in the text generation task (diffusionLM; instructCTG), we adopt it to assesses how smooth and natural the generated DNA sequence is. Fluency is defined as the perplexity of the generated DNA in a pretrained language model, with lower value indicating better fluency. We use the HyenaDNA pretrained model as the base model $F_{b}$ (hyenaDNA). Given the cross-entropy loss $\ell$, the fluency of a single DNA sequence $U=\{u_{1},\cdots,u_{n}\}$ is the exponential of the loss
@@
  • Diversity: In a specific cell type, a given protein can bind to many different DNA sequences. This metric measures the diversity of the generated sequences. Following discdiff, the diversity score is defined as
@@ 5.1 Promoter Generation
Dataset and task descriptions. This dataset contains 100,000 promoter sequences from GRCh38 (HG38) human reference genome. Each sequence is annotated with the transcription initiation profile, indicating the likelihood of transcription initiation activity at each base pair (fantom2014promoter). We use the same DNA sequences as in previous works (avdeyev2023dirichlet; stark2024dirichlet), specifically splitting 1,024 base pair long DNA sequences around each annotated transcription start site position (hon2017atlas). The task is to interpret the transcription initiation profile into the corresponding DNA sequences of the same length.
Metrics. Following prior works (avdeyev2023dirichlet; stark2024dirichlet), we evaluate the generation performance using the mean squared error (MSE) between the predicted regulatory activity of the generated and original DNA sequences. The regulatory activity is computed using the SEI framework (SEI), a published deep learning model that predicts 21,907 regulatory features for a given DNA sequence. For our primary evaluation, we extract promoter-related activity by focusing on the H3K4me3 chromatin mark. In addition, we report the Kolmogorov–Smirnov (KS) statistic between the distributions of regulatory activity values from the generated and original sequences, as proposed in (sarkar2024designing), to assess distributional alignment. We further evaluate the fluency of the generated sequences by HyenaDNA (nguyen2023hyenadna), as described in Section 4, to ensure that generated DNA is biologically plausible and structurally coherent.